Please help me figure out this problem... "14. Danny sold half of his comic books then bought 9 more. He now has 18 comic books. How many comic books did he begin with?"
Answer by TimothyLamb(4379)

x = original number of comics
---
x - x/2 + 9 = 18
(2/2)x - (1/2)x = 9
(1/2)x = 9
x = 9/(1/2)
x = 9*2
x = 18
---
answer:
x = original number of comic books = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php